OpenStudy (anonymous):

Which expression means 7 more than a number? dx7 or d+7

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle d \times 7 }\) or \(\large\color{black}{ \displaystyle d+7 }\) those are your options, right?

OpenStudy (solomonzelman):

(I thought u put a differential on there)

OpenStudy (solomonzelman):

anyway, which one do YOU think it is?

OpenStudy (anonymous):

i think it's +

OpenStudy (solomonzelman):

yes, correct

OpenStudy (solomonzelman):

\(\rm d\) is a number. And 7 more than \( \rm d\) means that you need to add 7 to \(\rm d\)